If the mass of the egg is 0.90 kg its initial speed is 2.5 m/s and the time to stop is 0.20 s find the average force on the egg?
Answer #1

Apply kinematic equation

v-u = at

a = v-u/t = 0-2.5/0.20 = -12.5 m/s^2

the average force on the egg is

F = 0.9(12.5) = 11.25 N